OpenStudy (anonymous):

Find the x and y intercepts and graph the equation plotting the intercepts -7x-8y=56

OpenStudy (mysesshou):

this one?

OpenStudy (anonymous):

this too lol but i was talking about my previous post problem

OpenStudy (mysesshou):

-7x-8y=56 Ok, x intercept is when y=0: -7x-0 =56 x=56/(-7) = -8 (-8,0) =x intercept then y intercept is when x=0 0-8y=56 y=56/(-8) y=-7 (0,-7) =y intercept plot those points and connect with a line

OpenStudy (mysesshou):

ok, send me the link of the one you want me to see. i didn't find it on the left side

OpenStudy (anonymous):

Write the slope-intercept form of the equation of the line shown in the graph. Please show all of your work. please see attachment

OpenStudy (mysesshou):

Ah. I think I started that one, but someone else had been working on it.

OpenStudy (mysesshou):

Ok.

OpenStudy (mysesshou):

So, it looks like the y intercept is approximately -1, where it crosses the y axis, but because the graph gives the blue dots showing specific points, then the y intercept is not necessarily exactly -1.

OpenStudy (mysesshou):

So, take the two blue points. (-6,-6)=(x1,y1) and (6,4)=(x2,y2) \[slope=m=\frac{y1-y2}{x1-x2}=\frac{-6-4}{-6-6}=\frac{-10}{-12}=\frac{5}{6}\]

OpenStudy (mysesshou):

Then there is that equation in the format (y-a)=m(x-b) ... i think.. Is that the right format?

OpenStudy (mysesshou):

So, seeing the equation in Wiki, they use a and b as the point#1 (y-(-6)) = (5/6)*(x-(-6)) y+6 = (5/6)*(x+6) y+6 = (5/6)*x+5 subtract 6 both sides y = (5/6)*x-1

OpenStudy (mysesshou):

Ok, so it turns out that the -1 does cross exactly at -1 on the y axis. The above method is for when you're not 100% sure. If you were sure, you could use the format y=mx+b put b=-1 and m=(5/6) directly giving y = (5/6)x-1

OpenStudy (mysesshou):

you can also count the slope. like start at (0,-1), and count 5 squares up, 6 squares to the right, and you land on that blue dot.

OpenStudy (mysesshou):

Is this ok?

OpenStudy (mysesshou):

message me if you need more info :)

OpenStudy (mysesshou):

what did the other person tell you?

OpenStudy (mysesshou):

I'm not sure I solved your issue... what was your question? so you had some of it, but the graph didn't turn out properly?

OpenStudy (anonymous):

okay i will the other person said the same

OpenStudy (mysesshou):

what was the confusing part?

OpenStudy (anonymous):

all of it

OpenStudy (mysesshou):

Ah. I'm not sure how to do any better. O_O

OpenStudy (anonymous):

its okay we working it out lol

OpenStudy (mysesshou):

phone call. just a bit